The question is as follows:
You have at least $30 in change in your drawer, consisting of
dimes and quarters. Write an inequality that shows the different number of coins in your drawer.
HELP!!!!

Answer by narayaba(40)

Let x be the number of dimes
Let y be the number of nickels
A dime is 10 cents equivalent to $(10/100) = $0.1
A nickel is 5 cents equivalent to $(5/100) = $0.05
The change in the drawer is atleast $30 which means the change in the drawer can be either $30 or more than $30
the total amount in dollars from the combination of dime and nickls
is 0.1*x + 0.05*y. This amount must be greater than or equal to 30.
0.1*x + 0.05*y >= 30